# inkscapelatex=false

### v1

@@ 4. Tokens evolving on an hemisphere
since $\alpha_{ij}^{\eta}(t,\mathbf{y})>0$ for each $1\leq i,j\leq\ell$ and $1\leq\eta\leq h$. Therefore, $\dot{\rm y}_{i}$ points to the interior of $\mathcal{H}^{+}(v)$ for each $1\leq i\leq\ell$, which ensures that $\mathcal{H}^{+}(v)$ is forward invariant under (11).
On the other hand, an easy check shows that every $\mathbf{y}\in\mathcal{C}_{\ell}^{+}(v)$ is an equilibrium of (11). Let us define the following function:
@@
where $V_{i}(\mathbf{y})=1-v^{\top}\,{\rm y}_{i}$. Let $\mathbf{y}=({\rm y}_{1},\ldots,{\rm y}_{\ell}):[0,\infty[\to(\mathcal{E}_{W}^{n})^{\ell}$ be a solution of (4.1) with $\mathbf{y}(0)\in\mathcal{H}^{+}(v)^{\ell}$. Forward invariance ensures that $\mathbf{y}(t)\in\mathcal{H}^{+}(v)^{\ell}$ for each $t\geq 0$. Moreover, let $\mathcal{I}(t)=\{i\in\{1,\ldots,\ell\}\mid V(\mathbf{y}(t))=V_{i}(\mathbf{y}(t))\}$. Note that, for $i\in\mathcal{I}(t)$, we have that $v^{\top}\,{\rm y}_{i}\leq v^{\top}\,{\rm y}_{j}$ for each $1\leq j\leq\ell$, where we dropped the argument $t$ for simplicity. Given that ${\rm y}_{i}^{\top}\,W\,{\rm y}_{j}\leq 1$ (since all tokens lie on the ellipsoid) and $v^{\top}\,{\rm y}_{i}>0$ (by assumption), we get:
@@
From this and (12), the upper Dini derivative of $V(\mathbf{y}(t))$, $t\geq 0$, is given by:
@@
The equality holds if and only if $v^{\top}\,{\rm y}_{i}(t)=v^{\top}\,{\rm y}_{j}(t)$ and ${\rm y}_{i}(t)^{\top}\,W\,{\rm y}_{j}(t)=1$ for each $1\leq j\leq\ell$, i.e., if and only if $\mathbf{y}(t)\in\mathcal{C}_{\ell}^{+}(v)$. In addition, the minimum of bounded and uniformly continuous functions is bounded and uniformly continuous. Hence, Lemma 4.2 ensures that $\dot{V}^{+}$ is uniformly continuous on [0,[[0$\infty$[[ 0 , ∞ [. Therefore, $V$ is a strict Lyapunov function for $\mathcal{C}_{\ell}^{+}(v)$ on $\mathcal{H}^{+}(v)^{\ell}$, and we conclude by the Lyapunov-like theorem based on Barbalat’s lemma (cf. [31, Theorem 8.4]). ∎
@@
Let us restrict ourselves to the case where we have normalization to the sphere, i.e., $W=\mathbb{I}_{n+1}$. Wendel’s theorem (cf. [32, Eq. (1)]) gives the probability that $\ell$ tokens lie on the same hemisphere when distributed uniformly at random; namely:
@@
5. Auto-regressive self-attention with identity value matrix
The remainder of this paper addresses the auto-regressive case, that is, the case where the dynamics of each token only depends on itself and the previous tokens. This corresponds to the model (7) with the so-called auto-regressive self-attention matrix, i.e.:
$Z_{i}^{\eta}(t,\mathbf{y})=\sqrt{n+1}\,\sum_{j=1}^{i}\exp({\rm y}_{i}^{\top}\,P_{\eta}(t)\,{\rm y}_{j}),\quad\alpha_{ij}^{\eta}(t,\mathbf{y})=\left\{\begin{array}[]{ll}\displaystyle\frac{1}{Z_{i}^{\eta}(t,\mathbf{y})}\,\exp({\rm y}_{i}^{\top}\,P_{\eta}(t)\,{\rm y}_{j}),&i\geq j,\\ 0,&i<j.\end{array}\right.$
Given that the equations are decoupled, the solution of the $i$-th equation only depends on the first $i$-th initial conditions. Hence, given an initial condition $\mathbf{y}^{0}=({\rm y}_{1}^{0},\ldots,{\rm y}_{\ell}^{0})\in(\mathcal{E}_{W}^{n})^{\ell}$, we denote the solution of the $i$-th equation by ${\rm y}_{i}(\cdot,{\rm y}_{1}^{0},\ldots,{\rm y}_{i}^{0}):[0,\infty[\to\mathcal{E}_{W}^{n}$, and the solution of the system by:
$\mathbf{y}(\cdot,\mathbf{y}^{0})=({\rm y}_{1}(\cdot,{\rm y}_{1}^{0}),\ldots,{\rm y}_{\ell}^{0}(\cdot,{\rm y}_{1}^{0},\ldots,{\rm y}_{\ell}^{0})):[0,\infty[\to(\mathcal{E}_{W}^{n})^{\ell}.$
@@ 5. Auto-regressive self-attention with identity value matrix
(14) $\tilde{\alpha}_{ij}^{\eta}:[0,\infty[\to[0,\infty[,\quad t\mapsto\tilde{\alpha}_{ij}^{\eta}(t)=\alpha_{ij}^{\eta}(t,\mathbf{y}(t,\mathbf{y}^{0})),$
@@
In this section, we focus on the case where $U_{\eta}(t)=\mathbb{I}_{n+1}$ for each $t\geq 0$, and $P_{\eta}$ is bounded (as a function of time) for each head $1\leq\eta\leq h$. In addition, we choose $W=\mathbb{I}_{n+1}$, i.e., the tokens evolve on the sphere. Consequently, the model becomes:
(15) $\dot{\rm y}_{i}=T_{{\rm y}_{i}}\pi\cdot\left(\sum_{\eta=1}^{h}\sum_{j=1}^{i}\alpha_{ij}^{\eta}(t,\mathbf{y})\,{\rm y}_{j}\right)=\sum_{\eta=1}^{h}\sum_{j=1}^{i}\alpha_{ij}^{\eta}(t,\mathbf{y})\left({\rm y}_{j}-{\rm y}_{i}^{\top}\,{\rm y}_{j}\,{\rm y}_{i}\right),$
for each $1\leq i\leq\ell$, $t\geq 0$ and $\mathbf{y}=({\rm y}_{1},\ldots,{\rm y}_{\ell})\in(\mathbb{S}^{n})^{\ell}$. In order to show that the consensus set is asymptotically stable, we make use of the following lemma.
@@
Consider a point ${\rm y}^{0}\in\mathbb{S}^{n}$ and let $\tilde{\alpha}:[0,\infty[\to[c,\infty)$ be a continuously differentiable function for some $c\in\mathbb{R}^{+}$. The only equilibria ${\rm y}^{*}\in\mathbb{S}^{n}$ of the following differential equation:
(16) $\dot{\rm y}=\tilde{\alpha}(t)({\rm y}^{0}-{\rm y}^{\top}\,{\rm y}^{0}\,{\rm y}),\qquad t\geq 0,~{}{\rm y}\in\mathbb{S}^{n},$
@@
For the first part, note that the equation $\tilde{\alpha}(t)({\rm y}^{0}-({\rm y}^{*})^{\top}\,{\rm y}^{0}\,{\rm y}^{*})=0$ holds for each $t\geq 0$ if and only if $({\rm y}^{*})^{\top}\,{\rm y}^{0}\in\{-1,1\}$, where we have used that $\tilde{\alpha}$ is always positive and both ${\rm y}^{0}$ and ${\rm y}^{*}$ lie on the $n$-sphere, i.e., if and only if ${\rm y}^{*}=-{\rm y}^{0}$ or ${\rm y}^{*}={\rm y}^{0}$. For the second part, let ${\rm y}:[0,\infty[\to\mathbb{S}^{n}$ be a solution of (16) and define:
$a:[0,\infty[\to[-1,1],\quad t\mapsto a(t)=({\rm y}^{0})^{\top}\,{\rm y}(t).$
From (16), the dynamics of $a$ is easily seen to be given by:
@@
The equilibria of the previous ODE are $a^{*}=-1$ and $a^{*}=1$, which correspond to ${\rm y}^{*}=-{\rm y}^{0}$ and ${\rm y}^{*}={\rm y}^{0}$, respectively. To check that the former is unstable and the latter is asymptotically stable with domain of attraction $(-1,1]$, which corresponds to $\mathbb{S}^{n}-\{-{\rm y}^{0}\}$, we define:
@@
We have that $V(a)\geq 0$ for $-1\leq a\leq 1$ and the equality holds if and only if $a=1$. Moreover, its derivative is given by $V^{\prime}(a)=a^{2}-1$, whence:
@@
Since $V^{\prime}(a)\neq 0$ for each $-1<a<1$, we conclude that $V$ is a Lyapunov function for the equilibrium $a^{*}=1$ and its domain of attraction is $(-1,1]$. ∎
We now state and prove the main result of this section establishing asymptotic stability of the consensus set for almost all initial conditions.
@@
If there exists $b>0$ such that $\max_{1\leq\eta\leq h}\sup_{t\geq 0}\|P_{\eta}(t)\|<b$, then the consensus set:
@@
is asymptotically stable for the system (15) and the domain of attraction contains the following set:
@@
Let $\mathbf{y}^{0}=({\rm y}_{1}^{0},\ldots,{\rm y}_{\ell}^{0})\in(\mathbb{S}^{n})^{\ell}$. To begin with, note that $\dot{\rm y}_{1}=0$ and, thus, the solution of the first equation is constant:
${\rm y}_{1}(t,{\rm y}_{1}^{0})={\rm y}_{1}^{0},\qquad t\geq 0.$
By substituting this into the second equation, we may write $\dot{\rm y}_{2}=\sum_{\eta=1}^{h}\tilde{\alpha}_{21}^{\eta}(t)({\rm y}_{1}^{0}-{\rm y}_{2}^{\top}\,{\rm y}_{1}^{0}\,{\rm y}_{2})$. From Lemmas 4.1 and 5.1, we conclude that ${\rm y}_{2}^{*}={\rm y}_{1}^{0}$ is the only asymptotically stable equilibrium with domain of attraction $\mathbb{S}^{n}-\{-{\rm y}_{1}^{0}\}$. In particular, $\mathcal{C}_{2}$ is asymptotically stable for the subsystem of (15) given by the first two tokens, and the domain of attraction is $\mathcal{D}_{2}^{1}$.
We proceed by induction: given $2\leq i\leq\ell$, for each $12\leq j\leq i-$ 1 suppose that $\mathcal{C}_{j}$ is asymptotically stable for the subsystem of (15) given by the first $j$ tokens, and the domain of attraction contains $\mathcal{D}_{j}^{1}$.
@@
$e_{j}:[0,\infty[\to\mathbb{R}^{n+1},\quad t\mapsto e_{j}(t)={\rm y}_{j}(t,{\rm y}_{1}^{0},\dots,{\rm y}_{j}^{0})-{\rm y}_{1}^{0},\qquad 1\leq j\leq i-1.$
Although $e_{1}=0$, it will be convenient to consider the system given by $e=(e_{1},\ldots,e_{i-1})$. The dynamics of the $i$-th token may be written as:
@@
Analogous to the proof of Lemma 5.1, we define:
$a_{i}:[0,\infty[\to[-1,1],\quad t\mapsto a_{i}(t)=({\rm y}_{1}^{0})^{\top}\,{\rm y}_{i}(t,{\rm y}_{1}^{0},\ldots,{\rm y}_{i}^{0}).$
Its dynamics is readily obtained from (18):
@@
Note that $a_{i}^{*}=-1$ and $a_{i}^{*}=1$ are the only equilibria of the previous ODE when $e=(e_{2},\ldots,e_{i-1})=0$. Let us consider the function $V$ introduced in (17). For the dynamics (19), it satisfies:
@@
where we have used that there exist $c_{1},c_{2}>0$ such that $c_{1}\leq\tilde{\alpha}_{ij}^{\eta}(t)\leq c_{2}$ for each $t\geq 0$ (recall Lemma 4.1). Note that $V^{\prime}(1)=0$ and $V^{\prime}|_{(-1,1)}\neq 0$. As a result, the previous inequality, together with the fact that $V|_{(-1,1)}>0$ and $V(1)=0$, ensures that $V|_{(-1,1]}$ is an ISS-Lyapunov function for the equilibrium $a_{i}^{*}=1$ of (19) where the input is given by $e=(e_{1},\ldots,e_{i-1})$. We conclude that $a_{i}^{*}=-1$ is an unstable equilibrium whereas $a_{i}^{*}=1$ is ISS-stable on $(-1,1]$. For the system (18), this corresponds to ${\rm y}_{i}^{*}=-{\rm y}_{1}^{0}$ being unstable and ${\rm y}_{i}^{*}={\rm y}_{1}^{0}$ being ISS-stable on $\mathbb{S}^{n}-\{-{\rm y}_{1}^{0}\}$.
Lastly, if we regard the errors as functions of the initial conditions, i.e., $e(t)=e(t,{\rm y}_{1}^{0},\ldots,{\rm y}_{i-1}^{0})$, then the induction hypothesis ensures that $e^{*}=0$ is an asymptotically stable equilibrium and its domain of attraction contains $\mathcal{D}_{i-1}^{1}$. As a result, $(e^{*},{\rm y}_{i}^{*})=(0,{\rm y}_{1}^{0})$ is asymptotically stable for the cascade system $(e,{\rm y}_{i})$ and its domain of attraction contains $\mathcal{D}_{i}^{1}$ (cf. [33, Lemma 4.7]). ∎
Remark 5.1 (Regular value matrix with different choice of projection).
The results in this section can be applied to non-identity value matrices, i.e., $U\neq\mathbb{I}_{n+1}$. To that end, we need to substitute the projection $T\pi_{W}$ introduced in (3) by a different projection to the ellipsoid. More specifically, we restrict ourselves to the single-head case, $h=1$, assume that $U$ is regular and pick $W=U^{\top}\,U$, which is symmetric and positive-define by construction. Then, for each ${\rm y}\in\mathcal{E}_{W}^{n}$, we define a new projection as:
@@
for each $1\leq i\leq\ell$, $t\geq 0$ and $\mathbf{y}=({\rm y}_{1},\ldots,{\rm y}_{\ell})\in(\mathcal{E}_{W}^{n})^{\ell}$.
The change of coordinates ${\rm z}_{i}=U\,{\rm y}_{i}\in\mathbb{S}^{n}$ for each $1\leq i\leq\ell$ brings (20) into:

### v2

@@ 4. Full self-attention matrix
since $\alpha_{ij}^{\eta}(t,\mathbf{y})>0$ for each $1\leq i,j\leq\ell$ and $1\leq\eta\leq h$. Therefore, $\dot{\rm y}_{i}$ points to the interior of $\mathcal{H}^{+}(v)$ for each $1\leq i\leq\ell$, which ensures that $\mathcal{H}^{+}(v)^{\ell}$ is forward invariant for Eq.˜6 under ˜4.1.
On the other hand, an easy check shows that every $\mathbf{y}\in\mathcal{C}_{\ell}^{+}(v)$ is an equilibrium of Eq.˜6 under ˜4.1. Let us define the following function:
@@
where $V_{i}(\mathbf{y})=1-v^{\top}{\rm y}_{i}$. Let $\mathbf{y}=({\rm y}_{1},\ldots,{\rm y}_{\ell}):\mathbb{R}_{0}^{+}\to(\mathcal{E}_{W}^{n})^{\ell}$ be a solution of Theorem˜4.1 with $\mathbf{y}(0)\in\mathcal{H}^{+}(v)^{\ell}$. Forward invariance ensures that $\mathbf{y}(t)\in\mathcal{H}^{+}(v)^{\ell}$ for each $t\in\mathbb{R}_{0}^{+}$. Moreover, let $\mathcal{I}(t)=\{i\in\{1,\ldots,\ell\}\mid V(\mathbf{y}(t))=V_{i}(\mathbf{y}(t))\}$. Note that, for $i\in\mathcal{I}(t)$, we have that $v^{\top}{\rm y}_{i}\leq v^{\top}{\rm y}_{j}$ for each $1\leq j\leq\ell$, where we dropped the argument $t$ for simplicity. Given that ${\rm y}_{i}^{\top}W{\rm y}_{j}\leq 1$ (since all tokens lie on the ellipsoid) and $v^{\top}{\rm y}_{i}>0$ (by assumption), we get:
@@
From this and Eq.˜9, the upper Dini derivative of $V(\mathbf{y}(t))$, $t\in\mathbb{R}_{0}^{+}$, is given by:
@@
The equality holds if and only if $v^{\top}{\rm y}_{i}(t)=v^{\top}{\rm y}_{j}(t)$ and ${\rm y}_{i}(t)^{\top}W{\rm y}_{j}(t)=1$ for each $1\leq j\leq\ell$, i.e., if and only if $\mathbf{y}(t)\in\mathcal{C}_{\ell}^{+}(v)$. In addition, the minimum of bounded and uniformly continuous functions is bounded and uniformly continuous. Hence, Lemma˜4.2 ensures that $\dot{V}^{+}$ is uniformly continuous on $\mathbb{R}_{0}^{+}$. Therefore, $V$ is a strict Lyapunov function for $\mathcal{C}_{\ell}^{+}(v)$ on $\mathcal{H}^{+}(v)^{\ell}$, and we conclude by the Lyapunov-like theorem based on Barbalat’s lemma (cf. Theorem 8.4 of [32]). ∎
Remark 4.1 (Autonomous systems).
For the case where the matrices $\{P_{\eta}\mid 1\leq\eta\leq h\}$ are time-independent, attractivity to the consensus set in the previous theorem can be upgraded to asymptotic stability by applying Lasalle’s invariance principle instead of Barbalat’s lemma.
@@
Let us restrict ourselves to the case where we have normalization to the sphere, i.e., $W=\mathbb{I}_{n+1}$. Wendel’s theorem (cf. Eq. (1) of [33]) gives the probability that $\ell$ tokens lie on the same hemisphere when distributed uniformly at random; namely:
@@
5. Auto-regressive self-attention matrix
This section addresses the auto-regressive (also known as causal) case, that is, the case where the dynamics of each token only depends on itself and the previous tokens. This corresponds to the model Eq.˜6 with the so-called auto-regressive self-attention matrix, i.e.:
$\alpha_{ij}^{\eta}(t,\mathbf{y})=\left\{\begin{array}[]{ll}\displaystyle\frac{1}{Z_{i}^{\eta}(t,\mathbf{y})}\exp({\rm y}_{i}^{\top}P_{\eta}(t){\rm y}_{j}),&i\geq j,\\ 0,&i<j,\end{array}\right$.
$Z_{i}^{\eta}(t,\mathbf{y})=\sqrt{n+1}\sum_{j=1}^{i}\exp({\rm y}_{i}^{\top}P_{\eta}(t){\rm y}_{j}).$
Note that the equations are decoupled and, thus, the solution of the $i$-th equation only depends on the first $i$-th initial conditions. Hence, given an initial condition $\mathbf{y}^{0}=({\rm y}_{1}^{0},\ldots,{\rm y}_{\ell}^{0})\in(\mathcal{E}_{W}^{n})^{\ell}$, we denote the solution of the $i$-th equation by ${\rm y}_{i}(\cdot,{\rm y}_{1}^{0},\ldots,{\rm y}_{i}^{0}):\mathbb{R}_{0}^{+}\to\mathcal{E}_{W}^{n}$, and the solution of the system by:
$\mathbf{y}(\cdot,\mathbf{y}^{0})=({\rm y}_{1}(\cdot,{\rm y}_{1}^{0}),\ldots,{\rm y}_{\ell}^{0}(\cdot,{\rm y}_{1}^{0},\ldots,{\rm y}_{\ell}^{0})):\mathbb{R}_{0}^{+}\to(\mathcal{E}_{W}^{n})^{\ell}.$
@@ 5. Auto-regressive self-attention matrix
(11) $\tilde{\alpha}_{ij}^{\eta}:\mathbb{R}_{0}^{+}\to\mathbb{R}_{0}^{+},\quad t\mapsto\tilde{\alpha}_{ij}^{\eta}(t)=\alpha_{ij}^{\eta}(t,\mathbf{y}(t,\mathbf{y}^{0})),$
@@
5.1. Identity value matrix
Let us consider the case where $W=\mathbb{I}_{n+1}$, i.e., the tokens evolve on the sphere.
Hypothesis 5.1.
The model Eq.˜6 is auto-regressive, $W=\mathbb{I}_{n+1}$ and, for each head $1\leq\eta\leq h$, we have:
  1. (1)
    $U_{\eta}(t)=\mathbb{I}_{n+1}$, and
  2. (2)
    $P_{\eta}(t)$ is bounded, i.e., $\sup_{t\in\mathbb{R}_{0}^{+}}\|P_{\eta}(t)\|<\infty$.
It is straightforward that, under the previous conditions, $\dot{\rm y}_{1}=0$ and, thus, the first token remains fixed: ${\rm y}_{1}(t)={\rm y}_{1}^{0}$ for each $t\in\mathbb{R}_{0}^{+}$. For each other token ${\rm y}_{i}$, $2\leq i\leq\ell$, the inner product between ${\rm y}_{1}^{0}$ and ${\rm y}_{i}$, i.e., the (cosine of the) angle between them, provides a projection of the dynamics onto the real line. In other words, using this angle we construct a scalar differential equation governing the evolution of the projection of the token on the real line. It them becomes simple to construct a Lyapunov function for the second token, ensuring convergence to ${\rm y}_{1}^{0}$ for every initial condition except for ${\rm y}_{2}(0)=-{\rm y}_{1}^{0}$. For the remaining tokens, an input-to-state stability argument coupled with the triangular nature of the dynamics leads to the asymptotic stability of the consensus set for almost all initial conditions. More specifically, let us start by proving the following lemma.
@@ 5.1. Identity value matrix
Consider a point ${\rm y}^{0}\in\mathbb{S}^{n}$ and let $\tilde{\alpha}:\mathbb{R}_{0}^{+}\to[c,\infty[$ be a continuously differentiable function for some $c\in\mathbb{R}^{+}$. The only equilibria ${\rm y}^{*}\in\mathbb{S}^{n}$ of the following differential equation:
(12) $\dot{\rm y}=\tilde{\alpha}(t)({\rm y}^{0}-{\rm y}^{\top}{\rm y}^{0}{\rm y}),\qquad t\in\mathbb{R}_{0}^{+},~{\rm y}\in\mathbb{S}^{n},$
@@
For the first part, note that the equation $\tilde{\alpha}(t)({\rm y}^{0}-({\rm y}^{*})^{\top}{\rm y}^{0}{\rm y}^{*})=0$ holds for each $t\in\mathbb{R}_{0}^{+}$ if and only if $({\rm y}^{*})^{\top}{\rm y}^{0}\in\{-1,1\}$, where we have used that $\tilde{\alpha}$ is always positive and both ${\rm y}^{0}$ and $\rm y$$^{*}$ lie on the $n$-sphere, i.e., if and only if ${\rm y}^{*}=-{\rm y}^{0}$ or ${\rm y}^{*}={\rm y}^{0}$. For the second part, let ${\rm y}:\mathbb{R}_{0}^{+}\to\mathbb{S}^{n}$ be a solution of Eq.˜12 and define:
$a:\mathbb{R}_{0}^{+}\to[-1,1],\quad t\mapsto a(t)=({\rm y}^{0})^{\top}{\rm y}(t).$
From Eq.˜12, the dynamics of $a$ is easily seen to be given by:
@@
The equilibria of the previous ODE are $a^{*}=-1$ and $a^{*}=1$, which correspond to ${\rm y}^{*}=-{\rm y}^{0}$ and ${\rm y}^{*}={\rm y}^{0}$, respectively. To check that the former is unstable and the latter is asymptotically stable with domain of attraction ]1,1]]-1,1], which corresponds to $\mathbb{S}^{n}-\{-{\rm y}^{0}\}$, we define:
@@
We have that $V(a)\in\mathbb{R}_{0}^{+}$ for $1\leq a\leq 1$ and $V(a)=0$ if and only if $a=1$. Moreover, its derivative is given by $V^{\prime}(a)=a^{2}-1$, whence:
@@
Since $V^{\prime}(a)\neq 0$ for each 1<a<1-1<a<1, we conclude that $V$ is a Lyapunov function for the equilibrium $a^{*}=1$ and its domain of attraction is ]1,1]]-1,1]. ∎
@@
If ˜5.1 holds, then the consensus set:
@@
is asymptotically stable for the system Eq.˜6 and the domain of attraction contains the following set:
@@
Let $\mathbf{y}^{0}=({\rm y}_{1}^{0},\ldots,{\rm y}_{\ell}^{0})\in(\mathbb{S}^{n})^{\ell}$. To begin with, note that $\dot{\rm y}_{1}=0$ and, thus, the solution of the first equation is constant, i.e., ${\rm y}_{1}(t,{\rm y}_{1}^{0})={\rm y}_{1}^{0}$ for each $t\in\mathbb{R}_{0}^{+}$. By substituting this into the second equation, we may write $\dot{\rm y}_{2}=\sum_{\eta=1}^{h}\tilde{\alpha}_{21}^{\eta}(t)({\rm y}_{1}^{0}-{\rm y}_{2}^{\top}{\rm y}_{1}^{0}{\rm y}_{2})$. From Lemmas˜4.1 and 5.1, we conclude that ${\rm y}_{2}^{*}={\rm y}_{1}^{0}$ is the only asymptotically stable equilibrium with domain of attraction $\mathbb{S}^{n}-\{-{\rm y}_{1}^{0}\}$. In particular, $\mathcal{C}_{2}$ is asymptotically stable for the subsystem of Eq.˜6 under ˜5.1 given by the first two tokens, and the domain of attraction is $\mathcal{D}_{2}^{1}$.
We proceed by induction: given $2\leq i\leq\ell$, for each $2\leq j\leq i-1$ suppose that $\mathcal{C}_{j}$ is asymptotically stable for the subsystem of Eq.˜6 under ˜5.1 given by the first $j$ tokens, and the domain of attraction contains $\mathcal{D}_{j}^{1}$.
@@
$e_{j}:\mathbb{R}_{0}^{+}\to\mathbb{R}^{n+1},\quad t\mapsto e_{j}(t)={\rm y}_{j}(t,{\rm y}_{1}^{0},\dots,{\rm y}_{j}^{0})-{\rm y}_{1}^{0},\qquad 1\leq j\leq i-1.$
Although $e_{1}=0$, it will be convenient to consider $e=(e_{1},\ldots,e_{i-1})$. The dynamics of the $i$-th token may be written as:
@@
Analogous to the proof of Lemma˜5.1, we define:
$a_{i}:\mathbb{R}_{0}^{+}\to[-1,1],\quad t\mapsto a_{i}(t)=({\rm y}_{1}^{0})^{\top}{\rm y}_{i}(t,{\rm y}_{1}^{0},\ldots,{\rm y}_{i}^{0}).$
Its dynamics is readily obtained from Eq.˜14:
@@
Note that $a_{i}^{*}=-1$ and $a_{i}^{*}=1$ are the only equilibria of the previous ODE when $e=(e_{2},\ldots,e_{i-1})=0$. Let us consider the function $V$ introduced in Eq.˜13. For the dynamics Eq.˜15, it satisfies:
@@
where we have used that there exist $c_{1},c_{2}>0$ such that $c_{1}\leq\tilde{\alpha}_{ij}^{\eta}(t)\leq c_{2}$ for each $t\in\mathbb{R}_{0}^{+}$ (recall Lemma˜4.1). Note that $V^{\prime}(1)=0$ and $V^{\prime}|_{]-1,1[}\neq 0$. As a result, the previous inequality, together with the fact that $V|_{]-1,1[}>0$ and $V(1)=0$, ensures that $V|_{]-1,1]}$ is an ISS-Lyapunov function for the equilibrium $a_{i}^{*}=1$ of Eq.˜15 where the input is given by $e=(e_{1},\ldots,e_{i-1})$. We conclude that $a_{i}^{*}=-1$ is an unstable equilibrium whereas $a_{i}^{*}=1$ is ISS-stable on ]1,1]]-1,1]. For the system Eq.˜14, this corresponds to ${\rm y}_{i}^{*}=-{\rm y}_{1}^{0}$ being unstable and ${\rm y}_{i}^{*}={\rm y}_{1}^{0}$ being ISS-stable on $\mathbb{S}^{n}-\{-{\rm y}_{1}^{0}\}$.
Lastly, if we regard the errors as functions of the initial conditions, i.e., $e(t)=e(t,{\rm y}_{1}^{0},\ldots,{\rm y}_{i-1}^{0})$, then the induction hypothesis ensures that $e^{*}=0$ is an asymptotically stable equilibrium and its domain of attraction contains $\mathcal{D}_{i-1}^{1}$. As a result, $(e^{*},{\rm y}_{i}^{*})=(0,{\rm y}_{1}^{0})$ is asymptotically stable for the cascade system $(e,{\rm y}_{i})$ and its domain of attraction contains $\mathcal{D}_{i}^{1}$ (cf. Lemma 4.7 of [34]). ∎
Remark 5.1 (Closest result available in the literature).
Remark 5.2 (Invertible value matrix with different choice of projection).
The results in this section can be applied to non-identity value matrices, i.e., $U\neq\mathbb{I}_{n+1}$. To that end, we need to substitute the projection $T\pi_{W}$ introduced in Eq.˜2 by a different projection to the ellipsoid. More specifically, we restrict ourselves to the single-head case, $h=1$, assume that $U$ is invertible and pick $W=U^{\top}U$, which is symmetric and positive-define by construction. Then, for each ${\rm y}\in\mathcal{E}_{W}^{n}$, we define a new projection as:
@@
for each $1\leq i\leq\ell$, $t\in\mathbb{R}_{0}^{+}$ and $\mathbf{y}=({\rm y}_{1},\ldots,{\rm y}_{\ell})\in(\mathcal{E}_{W}^{n})^{\ell}$.
The change of coordinates ${\rm z}_{i}=U{\rm y}_{i}\in\mathbb{S}^{n}$ for each $1\leq i\leq\ell$ brings Eq.˜16 into: